OpenStudy (anonymous):

A 12 ft. flag pole is mounted on top of a school building. If the top of the flagpole forms a 39° angle with the ground 50 ft. from the base of the building, about how tall is the school building? A. 27 ft. B. 28 ft. C. 19 ft. D. 40 ft.

OpenStudy (anonymous):

Try drawing a picture of what this is describing... You will use trig ratios to solve.

OpenStudy (anonymous):

@sashley0915 i dont know the trig ratios in order to solve this do you think you can help me out on this one?

OpenStudy (anonymous):

Look at the attached pic.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\sin(39)=\frac{x+12}{50}\]solve for x

OpenStudy (anonymous):

Then use SOH-CAH-TOA. sin(angle)=opposite/hypotenuse cos(angle)=adjacent/hypotenuse tan(angle)=opposite/hypotenuse. You will have decide which one to use based on the given information.

OpenStudy (anonymous):

@sashley0915 ugh man this is still not enough help for me based on your knowledge what do you think is the correct answer choice

OpenStudy (anonymous):

@satellite73 you do not want to use sin here because you are given the opposite and adjacent sides, you want to use the tan function.

OpenStudy (anonymous):

i want the tangent funtion and i clearly made a mistake

OpenStudy (anonymous):

@inflatablesteez Are you still having trouble understanding what happened here? satelitte's equation was correct except instead of sin use the tan function.

OpenStudy (anonymous):

should be \[\tan(39)=\frac{x+12}{50}\] \[50\tan (39)=x+12\] \[x=50\tan(39)-12\] and then a calculator

OpenStudy (anonymous):

@satellite73 so would the correct answer choice be B.28?

OpenStudy (anonymous):

@Hero help me

OpenStudy (anonymous):

@HyperChemist do you know the answer?

OpenStudy (anonymous):

@inflatablesteez I agree with the answer you calculated. (B)

OpenStudy (anonymous):

i typed the info in the calculator and it gave me 24 with decimals soo im going with 24 thank you to everyone who helped